# Large Stepsize Gradient Descent for Logistic Loss: Non-Monotonicity of the Loss Improves Optimization Efficiency

### v1

@@ Appendix D A Uniform-Convergence Generalization Bound
Proof of Proposition 15.
Consider a hypothesis class of hyperplanes
@@
We complete the proof by noting that $\hat{\mathbf{w}}$ can be any EMR. ∎
Appendix E Proof of Theorem 4
@@
The following lemma for SGD is an analogy of Lemma 7 for GD.
Lemma 16.
@@
Proof of Lemma 16.
The proof repeats the arguments in Lemma 7. By the SGD update rule, we have
@@
dividing both sides by $t$ completes our proof. ∎
Similarly, the following lemma for SGD is an analogy of Lemma 8 for GD.
Lemma 17 (Parameter and risk bounds).
@@
Proof of Lemma 17.
In Lemma 16, choose
@@
which verifies the first claim. For the second claim, we use Lemma 16 for $s\leq t$ and get
@@
We complete the proof. ∎
The next lemma provides a population risk bound. The main idea is to use martingale concentration tools to convert the “regret” bound in Lemma 17 to a population risk bound.
Lemma 18 (Population risk bound).
@@
Proof of Lemma 18.
Based on the uniform upper bound on the norm of $(\mathbf{w}_{k})_{k\leq t}$ in Lemma 17, we can show the following uniform upper bound on risk:
@@
Using Lemma 17 and the definition of $M$, we get
@@
which concludes our analysis. ∎
@@
Lemma 19 (Population error bound).
@@
Proof of Lemma 19.
Repeating the perceptron argument as in Lemma 9 , we get
@@
Using the parameter norm upper bound in Lemma 17, we have
@@
We conclude our proof by noting that zero-one loss is bounded by $2G(\cdot)$. ∎
The above lemmas imply Theorem 4.
Proof of Theorem 4.
It follows from Lemmas 18 and 19. ∎
Appendix F Proof of Proposition 5
Proof of Proposition 5.
We now verify Proposition 5.
@@ Appendix F Proof of Proposition 5
We have verified all examples. ∎
@@
Proof of Lemma 20.
Recall that
@@
which completes the proof. ∎
Appendix G Proof of Theorem 6
@@
The proof of Theorem 6 breaks down into three parts. In Section G.1, we control the network properties at initialization; then in Section G.2, we bound the linearization error. These two parts use standard techniques in the NTK literature. Results in Sections G.1 and G.2 allow us to show that a good event, under which the network is well-behaved at initialization and the linearization error is small, happens with high probability provided there is a large enough network width. Finally, in Section G.3, we analyze the GD dynamics under general losses assuming that the good event holds.
It is worth noting that our analysis in Section G.3 is decoupled from Sections G.1 and G.2 except assuming the good event holds. Therefore, our results for GD under general losses in Section G.3 directly apply to a simpler case of training linear models with zero initialization, since the good event automatically holds in this case; we include this consequence explicitly as Theorem 35 in Appendix H.
G.1 Initilization Bounds
@@
Lemma 21 (NTK feature margin).
@@
Proof of Lemma 21.
This is Lemma 2.3 in (Ji and Telgarsky, 2019). Choose
@@
By Hoeffding’s inequality, we have with probability at least $1-\delta$,
@@
Lemma 22 (Initial function value).
@@
Proof of Lemma 22.
This is a variant of Lemma 2.4 in (Ji and Telgarsky, 2019). Note that $\mathbf{x}^{\top}\mathbf{w}_{0}^{(s)}\sim\mathcal{N}(0,\|\mathbf{x}\|^{2})$, $\|\mathbf{x}\|\leq 1$, and $\phi(\cdot)$ is $1$-Lipschitz, so $\phi(\mathbf{x}^{\top}\mathbf{w}_{0}^{(s)})$ is $1$-subGaussian. Therefore,
@@
This completes the proof. ∎
G.2 Linearization Error Bounds
@@
Lemma 23 (Linearization error).
@@
Proof of Lemma 23.
This is a variant of Lemma 4.1 in (Telgarsky, 2021). The proof is included for completeness.
@@
Then by Hoeffding’s inequality, with probability at least $1-\delta$, we have
@@
holds with probability at least $1-\delta$. ∎
Using Lemma 23, we can show that the linearization error is well-bounded when the network width is sufficiently large.
Lemma 24 (Linearization error conditions).
@@
Proof of Lemma 24.
These follow from our choice of $m$ and Lemma 23. ∎
G.3 Optimization Analysis under General Losses
@@
Let $\mathcal{E}$ be the intersection of the events in Lemmas 21, 22 and 24. Then we know $\mathcal{E}$ holds with probability at least $1-3\delta$ if the width $m$ is large enough as specified in Theorem 6. In what follows, we always operate under the good event $\mathcal{E}$. We denote the gradient potential by
@@
Our next lemma establishes both a uniform margin bound and a relative margin bound for parameters in a ball. The relative margin bound is crucial to get the right condition on the width $m$ and the ball radius $R$.
Lemma 25 (Margin in a parameter ball).
@@
Proof of Lemma 25.
The first claim is because
@@
$\geq\frac{9\gamma}{10}-\|\nabla f_{i}(\mathbf{w})-\nabla f_{i}(\mathbf{w}_{0})\|$ by Lemma 21
$\geq\frac{4\gamma}{5}.$ by Lemma 24
@@
$f_{i}(\mathbf{v})$ $=f_{i}(\mathbf{w})+\langle\nabla f_{i}(\mathbf{w}),\mathbf{v}-\mathbf{w}\rangle+\xi_{i}(\mathbf{v},\mathbf{w})$
$=\langle\nabla f_{i}(\mathbf{w}),\mathbf{w}_{0}\rangle+\xi_{i}(\mathbf{v},\mathbf{w})$ by the homogeneity of $f$
$\leq\langle\nabla f_{i}(\mathbf{w}),\mathbf{w}_{0}\rangle+\frac{\gamma}{10}\|\mathbf{w}-\mathbf{v}\|.$ by Lemma 24
@@
These complete the proof. ∎
@@
Lemma 26 (Gradient potential properties).
@@
Proof of Lemma 26.
The first claim is by
@@
$\geq-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}\big{(}y_{i}f_{i}(\mathbf{w})\big{)}\cdot\frac{4\gamma}{5}$ by Lemma 25
@@
We have completed the proof. ∎
In the next lemma, we redo Lemma 7 for nonlinear models with controlled linearization errors and general loss functions.
Lemma 27 (Split optimization).
@@
Proof of Lemma 27.
For $k<t$, we have
@@
$\leq\frac{1}{n}\sum_{i=1}^{n}\ell\big{(}\gamma(8\theta-\|\mathbf{w}_{k}-\mathbf{w}_{0}\|)/10+y_{i}f_{i}(\mathbf{w}_{0})\big{)}-L(\mathbf{w}_{t})$ by Lemma 25
@@
From Lemma 22 we have
@@
which completes the proof. ∎
The following lemma combines an induction argument with the arguments in Lemma 8.
Lemma 28 (Risk and parameter bounds in the EoS phase).
@@
Proof of Lemma 28.
We prove the claims by induction. For $t=0$, the two claims hold.
@@
So we can invoke Lemma 27. Recall that
@@
$\leq\frac{\rho(\gamma^{2}\eta t)}{\gamma^{2}\eta t},$ by Lemma 20
@@
These verify the claim for $t$. We have completed the proof. ∎
The following lemma is analogous to Lemma 9.
Lemma 29 (Gradient potential bound in the EoS phase).
@@
Proof of Lemma 29.
By Lemma 28, we know $\mathbf{w}_{t}\in\mathcal{B}$ for $t\leq T$. Then we use the perceptron argument (Novikoff, 1962),
@@
$\geq\langle\mathbf{w}_{t},\mathbf{w}_{*}\rangle-\frac{4\gamma\eta}{5n}\sum_{i=1}^{n}\ell^{\prime}\big{(}y_{i}f_{i}(\mathbf{w}_{t})\big{)}$ by Lemma 25
@@
Plugging in the parameter bound in Lemma 28 completes the proof. ∎
The following lemma is analogous to Lemma 10. However, our focus here is a nonlinear predictor that might not be twice differentiable. We address this issue using the self-boundedness of the loss in 3C and a slightly longer induction argument.
Lemma 30 (Stable phase).
@@
Proof of Lemma 30.
Note that $\mathbf{w}_{1},\dots,\mathbf{w}_{T}$ all belong to $\mathcal{B}$ by Lemma 28.
@@
$\leq\eta\cdot G(\mathbf{w}_{t})$ by Lemma 26
@@
Applying the gradient bound 9 and the local linearization error bound in Lemma 24,
@@
where the last inequality is because $\gamma\leq 1$. Taking the average over $i=1,\dots,n$, we get
@@
  • For $t=s$, using our assumption on $L(\mathbf{w}_{s})$ and Lemma 26, we have
@@
    From Lemma 26, we have
@@
We have proved all claims. ∎
@@
Proof of Lemma 31.
Since $\ell(\cdot)\geq 0$ by 3A, we have
@@
The following lemma is analogous to Lemma 11. Again, we use an additional induction argument to get a sharp width condition.
Lemma 32 (Convergence in the stable phase).
@@
Proof of Lemma 32.
The first upper bound on $L(\mathbf{w}_{s})$ enables Lemma 30 for $s$ onwards. Therefore we have for $k\geq 0$,
@@
The assumption that $L(\mathbf{w}_{s})\leq\ell(0)/n$ allows us to apply Lemma 31, so $y_{i}f_{i}(\mathbf{w}_{s})\geq 0$ and thus
@@
We complete our induction by using the monotonicity of $L(\mathbf{w}_{t})$ for $t\geq s$ by Lemma 30. ∎
@@
Lemma 33 (Phase transition time).
@@
Proof of Lemma 33.
Applying Lemma 28 with $t=\tau$, we have
@@
which implies that there exists $s\leq\tau$ for $L(\mathbf{w}_{s})$ satisfies the right hand side bound. ∎
@@
Lemma 34 (Phase transition time under exponential tail).
@@
Proof of Lemma 34.
Under 3D, we have,
@@
Applying Lemma 29 for $\tau$, we have
@@
We complete the proof by plugging in the upper bound on $G(\mathbf{w}_{s})$. ∎
The proof of Theorem 6 follows from the above lemmas.
Proof of Theorem 6.
It follows from Lemmas 28, 32, 33 and 34. ∎
Appendix H General Losses and Linear Model
We restate our Theorem 6 for a linear model under general loss functions.
Theorem 35 (General losses and linear model).

### v2

@@ Appendix E A Uniform-Convergence Generalization Bound
Proof E.2.
(of Proposition E.1 Consider a hypothesis class of hyperplanes
@@
We complete the proof by noting that $\hat{\mathbf{w}}$ can be any EMR.
Appendix F Proof of Theorem 3.1
@@
The following lemma for SGD is an analogy of Lemma B.1 for GD.
Lemma F.1.
@@
Proof F.2.
(of Lemma F.1 The proof repeats the arguments in Lemma B.1. By the SGD update rule, we have
@@
dividing both sides by $t$ completes our proof.
Similarly, the following lemma for SGD is an analogy of Lemma B.3 for GD.
Lemma F.3 (Parameter and risk bounds).
@@
Proof F.4.
(of Lemma F.3 In Lemma F.1, choose
@@
which verifies the first claim. For the second claim, we use Lemma F.1 for $s\leq t$ and get
@@
We complete the proof.
The next lemma provides a population risk bound. The main idea is to use martingale concentration tools to convert the “regret” bound in Lemma F.3 to a population risk bound.
Lemma F.5 (Population risk bound).
@@
Proof F.6.
(of Lemma F.5 Based on the uniform upper bound on the norm of $(\mathbf{w}_{k})_{k\leq t}$ in Lemma F.3, we can show the following uniform upper bound on risk:
@@
Using Lemma F.3 and the definition of $M$, we get
@@
which concludes our analysis.
@@
Lemma F.7 (Population error bound).
@@
Proof F.8.
(of Lemma F.7 Repeating the perceptron argument as in Lemma B.5 , we get
@@
Using the parameter norm upper bound in Lemma F.3, we have
@@
We conclude our proof by noting that zero-one loss is bounded by $2G(\cdot)$.
The above lemmas imply Theorem 3.1.
Proof F.9.
(of Theorem 3.1 It follows from Lemmas F.5 and F.7.
Appendix G Proof of Proposition 4.1
Proof G.1.
(of Proposition 4.1 We now verify Proposition 4.1.
@@
We have verified all examples.
@@
Proof G.3.
(of Lemma G.2 Recall that
@@
which completes the proof.
Appendix H Proof of Theorem 4.2
@@
The proof of Theorem 4.2 breaks down into three parts. In Section H.1, we control the network properties at the initialization; then in Section H.2, we bound the linearization error. These two parts use standard techniques in the NTK literature. Results in Sections H.1 and H.2 allow us to show that a good event, under which the network is well-behaved at initialization and the linearization error is small, happens with high probability provided there is a large enough network width. Finally, in Section H.3, we analyze the GD dynamics under general losses assuming that the good event holds.
It is worth noting that our analysis in Section H.3 is decoupled from Sections H.1 and H.2 except assuming the good event holds. Therefore, our results for GD under general losses in Section H.3 directly apply to a simpler case of training linear models with zero initialization, since the good event automatically holds in this case; we include this consequence explicitly as Theorem I.1 in Appendix I.
H.1 Initilization Bounds
@@
Lemma H.1 (NTK feature margin).
@@
Proof H.2.
(of Lemma H.1 This is Lemma 2.3 in (Ji and Telgarsky, 2019). Choose
@@
By Hoeffding inequality, we have with probability at least $1-\delta$,
@@
Lemma H.3 (Initial function value).
@@
Proof H.4.
(of Lemma H.3 This is a variant of Lemma 2.4 in (Ji and Telgarsky, 2019). Note that $\mathbf{x}^{\top}\mathbf{w}_{0}^{(s)}\sim\mathcal{N}(0,\|\mathbf{x}\|^{2})$, $\|\mathbf{x}\|\leq 1$, and $\phi(\cdot)$ is $1$-Lipschitz, so $\phi(\mathbf{x}^{\top}\mathbf{w}_{0}^{(s)})$ is $1$-subGaussian. Therefore,
@@
This completes the proof.
H.2 Linearization Error Bounds
@@
Lemma H.5 (Linearization error).
@@
Proof H.6.
(of Lemma H.5 This is a variant of Lemma 4.1 in (Telgarsky, 2021). The proof is included for completeness.
@@
Then by Hoeffding inequality, with probability at least $1-\delta$, we have
@@
holds with probability at least $1-\delta$.
Using Lemma H.5, we can show that the linearization error is well-bounded when the network width is sufficiently large.
Lemma H.7 (Linearization error conditions).
@@
Proof H.8.
(of Lemma H.7 These follow from our choice of $m$ and Lemma H.5.
H.3 Optimization Analysis under General Losses
@@
Let $\mathcal{E}$ be the intersection of the events in Lemmas H.1, H.3 and H.7. Then we know $\mathcal{E}$ holds with probability at least $1-3\delta$ if the width $m$ is large enough as specified in Theorem 4.2. In what follows, we always operate under the good event $\mathcal{E}$. We denote the gradient potential by
@@
Our next lemma establishes both a uniform margin bound and a relative margin bound for parameters in a ball. The relative margin bound is crucial to get the right condition on width $m$ and the ball radius $R$.
Lemma H.9 (Margin in a parameter ball).
@@
Proof H.10.
(of Lemma H.9 The first claim is because
@@
$\geq\frac{9\gamma}{10}-\|\nabla f_{i}(\mathbf{w})-\nabla f_{i}(\mathbf{w}_{0})\|$ by Lemma H.1
$\geq\frac{4\gamma}{5}.$ by Lemma H.7
@@
$y_{i}f_{i}(\mathbf{v})$ $=y_{i}\big{(}f_{i}(\mathbf{w})+\langle\nabla f_{i}(\mathbf{w}),\mathbf{v}-\mathbf{w}\rangle+\xi_{i}(\mathbf{v},\mathbf{w})\big{)}$
$=y_{i}\big{(}\langle\nabla f_{i}(\mathbf{w}),\mathbf{v}\rangle+\xi_{i}(\mathbf{v},\mathbf{w})\big{)}$ by the homogeneity of $f$
$\leq y_{i}\langle\nabla f_{i}(\mathbf{w}),\mathbf{v}\rangle+\frac{\gamma}{10}\|\mathbf{w}-\mathbf{v}\|.$ by Lemma H.7
@@
These complete the proof.
@@
Lemma H.11 (Gradient potential properties).
@@
Proof H.12.
(of Lemma H.11 The first claim is by
@@
$\geq-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}\big{(}y_{i}f_{i}(\mathbf{w})\big{)}\cdot\frac{4\gamma}{5}$ by Lemma H.9
@@
We have completed the proof.
In the next lemma, we redo Lemma B.1 for nonlinear models with controlled linearization errors and general loss functions.
Lemma H.13 (Split optimization).
@@
Proof H.14.
(of Lemma H.13 For $k<t$, we have
@@
$\leq\frac{1}{n}\sum_{i=1}^{n}\ell\big{(}\gamma(8\theta-\|\mathbf{w}_{k}-\mathbf{w}_{0}\|)/10+y_{i}f_{i}(\mathbf{w}_{0})\big{)}-L(\mathbf{w}_{t})$ by Lemma H.9
@@
From Lemma H.3 we have
@@
which completes the proof.
The following lemma combines an induction argument with the arguments in Lemma B.3.
Lemma H.15 (Risk and parameter bounds in the EoS phase).
@@
Proof H.16.
(of Lemma H.15 We prove the claims by induction. For $t=0$, the two claims hold.
@@
So we can invoke Lemma H.13. Recall that
@@
$\leq\frac{\rho(\gamma^{2}\eta t)}{\gamma^{2}\eta t},$ by Lemma G.2
@@
These verify the claim for $t$. We have completed the proof.
The following lemma is analogous to Lemma B.5.
Lemma H.17 (Gradient potential bound in the EoS phase).
@@
Proof H.18.
(of Lemma H.17 By Lemma H.15, we know $\mathbf{w}_{t}\in\mathcal{B}$ for $t\leq T$. Then we use the perceptron argument (Novikoff, 1962),
@@
$\geq\langle\mathbf{w}_{t},\mathbf{w}_{*}\rangle-\frac{4\gamma\eta}{5n}\sum_{i=1}^{n}\ell^{\prime}\big{(}y_{i}f_{i}(\mathbf{w}_{t})\big{)}$ by Lemma H.9
@@
Plugging in the parameter bound in Lemma H.15 completes the proof.
The following lemma is analogous to Lemma B.7. However, our focus here is a nonlinear predictor that might not be twice differentiable. We address this issue using the self-boundedness of the loss in 3C and a slightly longer induction argument.
Lemma H.19 (Stable phase).
@@
Proof H.20.
(of Lemma H.19 Note that $\mathbf{w}_{1},\dots,\mathbf{w}_{T}$ all belong to $\mathcal{B}$ by Lemma H.15.
@@
$\leq\eta\cdot G(\mathbf{w}_{t})$ by Lemma H.11
@@
Applying the gradient bound 9 and the local linearization error bound in Lemma H.7,
@@
where the last inequality is because $\gamma\leq 1$. Taking average over $i=1,\dots,n$, we get
@@
  • For $t=s$, using our assumption on $L(\mathbf{w}_{s})$ and Lemma H.11, we have
@@
    From Lemma H.11, we have
@@
We have proved all claims.
@@
Proof H.22.
(of Lemma H.21 Since $\ell(\cdot)\geq 0$ by 3A, we have
@@
The following lemma is analogous to Lemma B.9. Again, we use an additional induction argument to get a sharp width condition.
Lemma H.23 (Convergence in the stable phase).
@@
Proof H.24.
(of Lemma H.23 The first upper bound on $L(\mathbf{w}_{s})$ enables Lemma H.19 for $s$ onwards. Therefore we have for $k\geq 0$,
@@
The assumption that $L(\mathbf{w}_{s})\leq\ell(0)/n$ allows us to apply Lemma H.21, so $y_{i}f_{i}(\mathbf{w}_{s})\geq 0$ and thus
@@
We complete our induction by using the monotonicity of $L(\mathbf{w}_{t})$ for $t\geq s$ by Lemma H.19.
@@
Lemma H.25 (Phase transition time).
@@
Proof H.26.
(of Lemma H.25 Applying Lemma H.15 with $t=\tau$, we have
@@
which implies that there exists $s\leq\tau$ for $L(\mathbf{w}_{s})$ satisfies the right hand side bound.
@@
Lemma H.27 (Phase transition time under exponential tail).
@@
Proof H.28.
(of Lemma H.27 Under 3D, we have,
@@
Applying Lemma H.17 for $\tau$, we have
@@
We complete the proof by plugging in the upper bound on $G(\mathbf{w}_{s})$.
The proof of Theorem 4.2 follows from the above lemmas.
Proof H.29.
(of Theorem 4.2 It follows from Lemmas H.15, H.23, H.25 and H.27.
Appendix I General Losses and Linear Model
We restate our Theorem 4.2 for a linear model under general loss functions.
Theorem I.1 (General losses and linear model).